OpenStudy (anonymous):

Sasha solved an equation. Her work is shown below. Explain Sasha's error. Then, show the correct way to solve. h divided by seven equals six. Add seven to both sides: h divided by seven plus seven equals six plus seven, so h equals thirteen.

OpenStudy (johnweldon1993):

\[h/7 = 6\] adding 7 to both sides will give h/7 + 7 = 13 that does not equal 6 + 7.....you need to multiply both sides by 7 h/7 times 7 gives h and 6 times 7 is 42 so h is 42

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

\[\frac{ h }{ 7 } = 6\] \[\frac{ h }{ 7 } + 7 = 6 + 7\] \[\frac{ h+49 }{ 7 } = 13\] \[h = 13*7 - 49\] = 42